OpenStudy (anonymous):

identify the maxiumum or minimum value of the function f(x)= (x+3)^2-11

OpenStudy (anonymous):

Find f´(x) and set f´(x)=0

OpenStudy (anonymous):

A) (-11,3) maximum B) (-3,-11) maximum C) (3,-11) minimum D) (-3,-11) minimum

OpenStudy (anonymous):

\[f(x)=(x+3)^2-11=x^2+6x-2\] what is f´(x)?

OpenStudy (anonymous):

3,-11?

OpenStudy (anonymous):

No... \[f'(x)=2x+6\]

OpenStudy (anonymous):

\[2x+6=0\] Find x

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

-3*

OpenStudy (anonymous):

Yes... Then find f(-3)

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

Yes.. :)

OpenStudy (anonymous):

idk if its a maximum or minimum

OpenStudy (anonymous):

You have to se its a parable.

OpenStudy (anonymous):

The function for a parable is f(x)=ax^2+bx+c

OpenStudy (anonymous):